Water from the Don Juan Pond in Antarctica is 4.40 M sodium chloride. If you have a 

340. mL sample of lake water, how many grams are in solution?


1
Expert's answer
2021-03-22T06:37:31-0400

Molarity = 4.40M

Volume of solution = 340 mL = 0.340L


Molarity = "\frac{no. of moles of solute }{volume of solution}"


4.40 M = "\frac{moles of solute }{0.340 L}"


moles of solute = 1.496 moles

molecular weight of sodium chloride = 58.44g

Grams in solution = 1.496×58.44 = 87.42g